OpenStudy (anonymous):

SIMPLIFY each expression. if not than write no. 3(5+6h)

OpenStudy (anonymous):

16 + 18h

OpenStudy (anonymous):

18h+15

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

since a(b+c) = (ab + bc) 3(5 + 6h) = 15 + 18h

OpenStudy (anonymous):

sorry i meant 15